# DiffHairCard: Auto Hair Card Extraction with Differentiable Rendering

### v1

@@ 6. Conclusion
We have presented a fully automated pipeline for converting strand-based hair models into efficient and visually compelling hair card representations. By leveraging a differentiable rendering framework, our method first clusters the hair strands to initialize card geometry, and then clusters the hair textures for sharing to reduce the memory cost. Finally, we conduct jointly optimization over both card geometry and textures. A key contribution of our approach is the introduction of 2D curve-based texture encoding, which offers a resolution-independent representation that effectively captures fine strand details while remaining compatible with differentiable rendering. Our method supports a wide range of hairstyles and lengths, and introduces mechanisms such as hair caps and crossed card generation to handle visually complex hair types like short and coily styles. Additionally, the ability to share textures across LoDs makes our approach well-suited for real-time applications where memory efficiency and performance are critical.
@@
While our pipeline demonstrates strong performance and visual fidelity across a range of hairstyles, several limitations remain. First, the optimization is conducted in multiple substeps rather than as a fully end-to-end differentiable pipeline, which may limit global consistency. Second, the loss functions primarily rely on visible attributes such as tangent alignment, depth, and mask coverage. However, for complex hairstyles, especially those with dense or layered structures, interior strand arrangements play an important visual role and are not explicitly enforced. Third, our method assumes a uniform hair color across the entire hairstyle, which may not generalize well to stylized or multi-colored hair. Finally, although our framework supports a wide range of hair types, coily hair remains a challenge due to its highly curved geometry, as shown in Fig. 13, which typically requires a large number of cards for faithful representation.

### v2

@@ 7. Conclusion
We have presented a fully automated pipeline for converting strand-based hair models into efficient and visually compelling hair card representations. By leveraging a differentiable rendering framework, our method first clusters the hair strands to initialize card geometry and then clusters the hair textures to share, thereby reducing memory cost. Finally, we conduct joint optimization over both card geometry and textures. A key contribution of our approach is the introduction of 2D curve-based texture encoding, which offers a resolution-independent representation that effectively captures fine strand details while remaining compatible with differentiable rendering. Our method supports a wide range of hairstyles and lengths, introducing mechanisms such as hair caps and crossed card generation to handle visually complex hair types, including short and coily styles. Additionally, the ability to share textures across LoDs makes our approach well-suited for real-time applications where memory efficiency and performance are critical.
@@
Although our pipeline demonstrates strong performance and visual fidelity across a range of hairstyles, several limitations remain. First, optimization is conducted in multiple substeps rather than as a fully end-to-end differentiable pipeline, which may limit global consistency. Given the already large number of variables to optimize (approximately 500K variables for the afro example), increasing the number of strand control points would indeed improve the representation capability of strands. However, this also enlarges the solution space, making optimization more prone to getting trapped in local minima. Second, the loss functions primarily rely on visible attributes such as tangent alignment, depth, and mask coverage. However, for complex hairstyles, especially those with dense or layered structures, interior strand arrangements play an important visual role and are not explicitly enforced. Besides, hair cards, as a low-cost representation, have an inherent limitation in that it is mostly for static and moderate animation in games. Under large deformation, it can cause occlusion artifacts. Third, our method assumes a uniform hair color throughout the entire hairstyle, which may not generalize well to stylized or multicolored hair. Finally, hair type with highly curved geometry, such as coily hair, remains a challenge. Theoretically, increasing the number of hair clusters and card geometries makes it easier to represent complex hairstyles. Ultimately, if we assign one card per strand, our pipeline would always succeed. However, our goal is to strike a balance between geometry and appearance. Fig. 16 shows that using too few hair cards, for example, only 20, fails to adequately represent the hair structure, even for relatively simple straight styles. As the number of cards increases, the overall fidelity of the hair representation improves. Similarly, Fig. 17 demonstrates that our method struggles to capture complex coily structures when both the number of hair cards and control points are limited. However, with a sufficiently large number of cards (e.g., $00×\times2$), even intricate coily patterns can be represented with fine detail.